### v1

@@ Abstract
We introduce and formalize the notion of resource augmentation for maximin share allocations — an idea that can be traced back to the seminal work of Budish [JPE 2011]. Specifically, given a fair division instance with $m$ goods and $n$ agents, we ask how many copies of the goods should be added in order to guarantee that each agent receives at least their original maximin share, or an approximation thereof. We establish a tight bound of $m/e$ copies for arbitrary monotone valuations. For additive valuations, we show that at most $\min\{n-2,\left\lfloor\frac{m}{3}\right\rfloor(1+o(1))\}$ copies suffice. For approximate-MMS in ordered instances, we give a tradeoff between the number of copies needed and the approximation guarantee. In particular, we prove that $\lfloor n/2\rfloor$ copies suffice to guarantee a $6/7$-approximation to the original MMS, and $\lfloor n/3\rfloor$ copies suffice for a $4/5$-approximation. Both results improve upon the best known approximation guarantees for additive valuations in the absence of copies.
@@ 1 Introduction
Clearly, the larger $d$ is, the easier it is to achieve the benchmark. The question then becomes, what is the smallest value of $d\geq n$ such that there is an allocation for $n$ agents so that each agent achieves the $1$-out-of-$d$ maximin share. This is also referred to as the ordinal approximation. In [31] the authors show that $1$-out-of-$\lceil 3n/2\rceil$ MMS allocations exist for any instance with additive valuations, later improved to $1$-out-of-$4\lceil n/3\rceil$ MMS allocations in [3], the current state of the art.
@@ 1.1 Our Contributions
Our main contribution is to introduce relaxations to the MMS by allowing resource augmentation, (i.e., allowing extra copies of items) and to obtain both positive and negative results in different settings. This relaxation was implicitly suggested by Budish in [19], where the number of items (slots in classrooms) tended to infinity. We ask:
@@
We consider both the total number of copies and the maximum number of copies of an item. We distinguish between distinct copies (only one extra copy of an item allowed), and when more copies of a single item are allowed. Much of this paper is focused on distinct copies —- up to one extra copy of an item. Our results are the following:
@@
$\alpha$-MMS Approximations with Copies for Ordered Instances (Additive Valuations)
We give new approximation guarantees for ordered instances, which improve upon the best-known approximations (without copies). Ordered instances, also called same ordered preferences (SOP) in [14], are such where all agents have additive valuations, and all agents have identical ranking of the items. We give points along a tradeoff between the guarantee of the $\alpha$-MMS approximation and the number of copies needed, we show how to achieve a $6/7$-MMS with $\lfloor n/2\rfloor$ distinct copies (Theorem 4) and how to achieve a $4/5$-MMS with $\lfloor n/3\rfloor$ distinct copies (Theorem 5).
@@ 1.2 Our Techniques
Our upper bounds for general monotone valuations use the probabilistic method. This is used for bounding the total number of copies as well as for bounding the maximal number of copies of a single item.
@@
  • We introduce bag filling with copies. Our bag filling techniques give full MMS allocations while duplicating some of the items. See Algorithm 2: $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$
@@
  • We combine bag filling and round robin approaches to achieve approximate MMS with copies for ordered instances. See Algorithm 4: $\mathsf{BagFill\mbox{-}RoundRobin}$.
@@ 1.4 Other Related Work
Fair division of indivisible goods is a vast area of research in recent years, a survey paper by Amanatidis et al, “Fair Division of Indivisible Goods: Recent Progress and Open Questions”, [6], includes 12 pages of references, and the authors stress that this is not an exhaustive survey and refer to yet other surveys (e.g., [34] with a more economic perspective) that focus on specific aspects of the problem.
@@ 2 Model and Preliminaries
As valuations can be scaled, we may assume, without loss of generality that the MMS value for every agent is one. For additive valuations, this implies that every agent values the grand bundle (all items) at $n$ or more.
@@
An allocation with copies is an allocation where goods may be duplicated, with different copies allocated to different agents. An agent may receive at most one copy of each good. An allocation with copies is a relaxation of an allocation where the two or more bundles may have a non-empty intersection.
@@
Note that, the definition above implies that no agent is allowed to receive more than a single copy of each good.
@@ 3.1 A Tight Bound of $m/e$ on the Total Number of Copies
Our lower bound is based on a combinatorial construction which considers $m=n^{n}$ goods positioned on the $n$-dimensional cube, so each good may be identified with $n$ coordinates, $(d_{1},\dots,d_{n})$. The valuations are such that each of the agent is almost “single-minded”, with agent $i$ only interested in bundles that contain all $n^{n-1}$ goods that share the same value in the $d_{i}$ coordinate. Those $n$ different bundles (one for each possible value of $d_{i}$), make up his MMS partition. A simple counting argument shows that the only way to satisfy all agents simultaneously is to copy $(n-1)^{n}$ goods, which is a $1/e$-fraction of the total number of goods, as $n$ tends to infinity.
@@
Consider $n$ agents and $n^{n}$ goods positioned on the $n$-dimensional cube. I.e., the set of goods is
@@
$\left|\bigcup_{i\in[n]}P_{i}^{j_{i}}\right|$ $=$ $\sum_{k=1}^{n}(-1)^{k+1}\binom{n}{k}n^{n-k}=\binom{n}{1}n^{n-1}-\left(\sum_{k=2}^{n}(-1)^{k}\binom{n}{k}n^{n-k}\right)$
$=$ $n^{n}-\left(n^{n}-n^{n}+\sum_{k=2}^{n}(-1)^{k}\binom{n}{k}n^{n-k}\right)=n^{n}-\left(\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}n^{n-k}\right)$
$=$ $n^{n}-(n-1)^{n},$
@@
$\mathbb{E}_{A}\left[|\cup_{i\in[n]}A_{i}|\right]$ $=$ $\sum_{g\in[m]}\Pr[g\in\cup_{i\in[n]}A_{i}]=\sum_{g\in[m]}1-\Pr[g\notin\cup_{i\in[n]}A_{i}]=\sum_{g\in[m]}1-\left(1-\frac{1}{n}\right)^{n}$
@@
so there must exists an allocation that assigns each agent one of her MMS bundles and requires no more than $m\left(1-\frac{1}{n}\right)^{n}$ extra copies. ∎
@@ 3.2 Bound on the Maximum Number of Copies of Any Good
Consider an allocation where agent $i$’s bundle, is randomly chosen as one of the sets in $i$’s MMS partition. We view this process as throwing throwing $n$ balls (agents) into $n$ bins (the index in an MMS partition). Fix some good $g\in M$, it belongs in to some MMS set for all the agents, possibly different indices, but we can rename the sets and assume without loss of generality that $g$ is always in the set of index 1 in the MMS partitions of all agents. The event in which $g$ is copied at least $k$ times for the allocation $A=(A_{1},\dots,A_{n})$, is exactly the event in which the bin corresponding to index $1$ has at least $k$ balls. For $k=\frac{3\ln m}{\ln\ln m}$, the probability of having a bin with more than $k$ balls is at most $\frac{1}{m}$. ∎
@@
Fix an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$. For any agent $i$, let $P^{i}=(P^{i}_{1},\dots,P^{i}_{n})$ be his MMS partition. Consider the algorithm $\mathcal{A}$ which does the following (at most) $3m\beta$ times:
@@ 4.1 Full MMS with $n-2$ Copies
For any fair division instance with additive valuations, at most $n-2$ distinct copies suffice to guarantee a full MMS allocation.
@@
$\mathsf{Match\mbox{-}n\mbox{-}Fill}$ first allocates items with value more than the MMS of an agent, so for every $i\in\tilde{N},j\in\tilde{M}$, $v_{ij}\leq 1$. Moreover, by Lemma 1, allocating a single item to a single agent is a valid reduction. Thus, after this step, we have that for every $i\in N$, $v_{i}(M)\geq|N|$. We show that after allocating bundles $\hat{P}$ to agents in $N$, for every unallocated agent $i\in\tilde{N}$, $v_{i}(\tilde{M})\geq|\tilde{N}|$.
@@
as desired. It follows that all the conditions required in order to apply Lemma 3 are fulfilled, and we can satisfy agents in $\tilde{N}$ using $|\tilde{N}|-1\leq|N|-2\leq n-2$ distinct copies (recall that at least one agent was already allocated before). Since $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ is the only place where we produce copies, this concludes the proof. ∎
@@
One can consider a similar question in the context of chore allocation, and ask “what is the maximum number of chores that can be assigned to agents such that no agent’s cost is larger than their maximin share”; that is, how many chores are needed to be set aside in order to produce a full MMS allocation. It turns out that an analogous process to $\mathsf{Match\mbox{-}n\mbox{-}Fill}$ in the context of chores outputs an allocation where at most $n-2$ items are disposed of, and each agent is allocated a bundle of cost no larger than their MMS value. More on this in Appendix A.
@@ 4.2 Full MMS with $\lfloor m/3\rfloor(1+o(1))$ Copies
$=$ $\left\lfloor\frac{\frac{\alpha}{1+\alpha}n+1}{\frac{1}{1+\alpha}n-1}m\right\rfloor=\left\lfloor\left(\alpha+\frac{1+\alpha}{\frac{1}{1+\alpha}n-1}\right)m\right\rfloor$
@@
5 Approximate-MMS with Copies: Ordered Instances
@@
In this section we study approximate MMS allocations for ordered instances while allowing duplication of goods. An instance is ordered if there exists a permutation of the goods $(g_{1},\ldots,g_{m})$ such that $v_{i}(g_{1})\geq\ldots\geq v_{i}(g_{m})$ for all $i\in N$ and also $v_{i}$ is additive for all $i\in N$.

### v2

@@ Abstract
We introduce and formalize the notion of resource augmentation for maximin share allocations — an idea that can be traced back to the seminal work of Budish [JPE 2011]. Specifically, given a fair division instance with $m$ goods and $n$ agents, we ask how many copies of the goods should be added in order to guarantee that each agent receives at least their original maximin share, or an approximation thereof. We establish a tight bound of $m/e$ copies for arbitrary monotone valuations. For additive valuations, we show that at most $\min\{n-2,\left\lfloor\frac{m}{3}\right\rfloor(1+o(1))\}$ copies suffice. For approximate-MMS in instances with additive valuations, we give a tradeoff between the number of copies needed and the approximation guarantee. In particular, we prove that $\lfloor n/2\rfloor$ copies suffice to guarantee a $6/7$-approximation to the original MMS, and $\lfloor n/3\rfloor$ copies suffice for a $4/5$-approximation. Both results improve upon the best known approximation guarantees for additive valuations in the absence of copies.
@@ 1 Introduction
Clearly, the larger $d$ is, the easier it is to achieve the benchmark. The question then becomes, what is the smallest value of $d\geq n$ such that there is an allocation for $n$ agents so that each agent achieves the $1$-out-of-$d$ maximin share. This is also referred to as the ordinal approximation. In [33] the authors show that $1$-out-of-$\lceil 3n/2\rceil$ MMS allocations exist for any instance with additive valuations, later improved to $1$-out-of-$4\lceil n/3\rceil$ MMS allocations in [4], the current state of the art. For all we know, even $1$-out-of-$(n+1)$ MMS allocations may exist.
@@ 1.1 Our Contributions
Our main contribution is to introduce relaxations to the MMS by allowing resource augmentation, (i.e., allowing extra copies of items) and to obtain both positive and negative results in different settings. This relaxation was implicitly suggested by Budish in [21], in the context of course allocation (where the items are seats in a classroom), and it may also be applicable to other settings in which adding copies of an item is possible (albeit with some restrictions), e.g., assigning private tutors to students, caregivers to elderly individuals, or distributing any form of hourly-based service. We ask:
@@
We consider both the total number of copies and the maximum number of copies of an item. We distinguish between distinct copies (only one extra copy of an item allowed), and when more copies of a single item are allowed. Much of this paper is focused on distinct copies — up to one extra copy of an item. Our results are the following:
@@
$\alpha$-MMS Approximations with Copies for Additive Valuations.
We give new approximation guarantees for additive instances, which improve upon the best-known approximations without copies. We give points along a tradeoff between the guarantee of the $\alpha$-MMS approximation and the number of copies needed. We show how to achieve a $6/7$-MMS with $\lfloor n/2\rfloor$ distinct copies (Theorem 4), and how to achieve a $4/5$-MMS with $\lfloor n/3\rfloor$ distinct copies (Theorem 5).
Relation to 1-out-of-$d$ MMS.
As mentioned, we show 1-out-of-$(1+\alpha)n$ MMS implies that exact MMS can be achieved with $\alpha m(1+o(1))$ distinct copies (see Lemma 5). In the other direction, showing that $n/\beta$ copies are sufficient for achieving exact MMS, immediately implies that 1-out-of-$(1-\beta)n$ MMS allocations exist. This is because at least $(1-\beta)n$ agents do not receive any copies, and obtain their MMS value. In particular, for additive valuations, achieving MMS with less than $n/4$ copies improves upon the currently best-known bound for 1-out-of-$d$.
@@ 1.2 Our Techniques
Our upper bounds for general monotone valuations use the probabilistic method. This is used for bounding the total number of copies as well as for bounding the maximum number of copies of a single item.
@@
  • We introduce bag filling with copies. Our bag filling techniques give full MMS allocations while duplicating some of the items. See Algorithm 2: $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$.
@@
  • While it was shown by [16] that considering agents with the same order of valuations over the items is without loss of generality when studying MMS, their proof does not carry over to the setting with copies. Nevertheless, using an augmenting path argument, we prove that for a restricted class of allocations, in particular those produced by our algorithms, such a reduction applies (see Section 5). Consequently, our algorithms for achieving approximate MMS assume, without loss of generality, that the instance is ordered. It is an intriguing open problem whether such a reduction applies in general.
  • We combine bag filling and round robin approaches to achieve approximate MMS allocations with copies for ordered instances. See Algorithm 4: $\mathsf{BagFill\mbox{-}RoundRobin}$.
@@ 1.4 Other Related Work
Fair division of indivisible goods is a vast area of research in recent years, a survey paper by Amanatidis et al., “Fair Division of Indivisible Goods: Recent Progress and Open Questions”, [7], includes 12 pages of references, and the authors stress that this is not an exhaustive survey and refer to yet other surveys (e.g., [36] with a more economic perspective) that focus on specific aspects of the problem.
@@
The work of [2] studies EF1 and Pareto optimal allocations of chores under chore duplication, whereas our work focuses on chore deletion. In a similar spirit to our resource augmentation approach, [41] and [15] investigate the computational complexity of achieving envy-free and proportional allocations when a limited number of goods can be divided among multiple agents.
@@ 2 Model and Preliminaries
Without loss of generality, we remove all agents whose MMS value is 00. These agents get their full MMS with empty bundles. As valuations can be scaled, we may assume, without loss of generality that the MMS value for every agent is one. For additive valuations, this implies that every agent values the grand bundle (all items) at $n$ or more.
@@
An allocation with copies is an allocation where goods may be duplicated, with different instances of a same good allocated to different agents. An agent may receive at most one instance of each good. An allocation with copies is a relaxation of an allocation where the two or more bundles may have a non-empty intersection.
@@
Note that, the definition above implies that no agent is allowed to receive more than a single instance of each good.
@@ 3.1 A Tight Bound of $m/e$ on the Total Number of Copies
Our lower bound is based on a combinatorial construction which considers $m=n^{n}$ goods positioned on the $n$-dimensional cube, so each good may be identified with $n$ coordinates, $(d_{1},\dots,d_{n})$. The valuations are such that each of the agent is almost “single-minded”, with agent $i$ only interested in bundles that contain all $n^{n-1}$ goods that share the same value in the $d_{i}$ coordinate. Those $n$ different bundles (one for each possible value of $d_{i}$), make up her MMS partition. A simple counting argument shows that the only way to satisfy all agents simultaneously is to copy $(n-1)^{n}$ goods, which is a $1/e$-fraction of the total number of goods, as $n$ tends to infinity.
@@
Consider $n$ agents and $n^{n}$ goods positioned on the $n$-dimensional cube, i.e., the set of goods is
@@
$\left|\bigcup_{i\in[n]}P_{i}^{j_{i}}\right|$ $=$ $\sum_{k=1}^{n}(-1)^{k+1}\binom{n}{k}n^{n-k}=n^{n}-n^{n}-\left(\sum_{k=1}^{n}(-1)^{k}\binom{n}{k}n^{n-k}\right)$
$=$ $n^{n}-\left(\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}n^{n-k}\right)=n^{n}-(n-1)^{n},$
@@
$\mathbb{E}_{A}\left[|\cup_{i\in[n]}A_{i}|\right]$ $=$ $\sum_{g\in[m]}\Pr[g\in\cup_{i\in[n]}A_{i}]=\sum_{g\in[m]}(1-\Pr[g\notin\cup_{i\in[n]}A_{i}])=\sum_{g\in[m]}\left(1-\left(1-\frac{1}{n}\right)^{n}\right)$
@@
so there must exists an allocation that assigns each agent one of her MMS bundles and requires no more than $m\cdot\left(1-\frac{1}{n}\right)^{n}$ extra copies. ∎
@@ 3.2 Bound on the Maximum Number of Copies of Any Good
Consider an allocation where agent $i$’s bundle, is randomly chosen as one of the sets in $i$’s MMS partition. We view this process as throwing $n$ balls (agents) into $n$ bins (the index in an MMS partition). Fix some good $g\in M$, it belongs to some MMS set for all the agents, possibly different indices, but we can rename the sets and assume without loss of generality that $g$ is always in the set of index 1 in the MMS partitions of all agents. The event in which $g$ is copied at least $k$ times for the allocation $A=(A_{1},\dots,A_{n})$, is exactly the event in which the bin corresponding to index $1$ has at least $k$ balls. For $k=\frac{3\ln m}{\ln\ln m}$, the probability of having a bin with more than $k$ balls is at most $\frac{1}{m}$. ∎
@@
Fix an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$. For any agent $i$, let $P^{i}=(P^{i}_{1},\dots,P^{i}_{n})$ be her MMS partition. Consider the algorithm $\mathcal{A}$ which does the following (at most) $3m\beta$ times:
@@ 4.1 Full MMS with $n-2$ Copies
For any fair division instance with additive valuations, Algorithm $\mathsf{Match\mbox{-}n\mbox{-}Fill}$ guarantees a full MMS allocation for all the agents with at most $n-2$ distinct copies.
@@
$\mathsf{Match\mbox{-}n\mbox{-}Fill}$ first allocates items with value more than the MMS of an agent, so for every $i\in\tilde{N},j\in\tilde{M}$, $v_{ij}<1$. Moreover, by Lemma 1, allocating a single item to a single agent is a valid reduction. Thus, after this step, we have that for every $i\in N$, $v_{i}(M)\geq|N|$. We show that after allocating bundles $\hat{P}$ to agents in $N$, for every unallocated agent $i\in\tilde{N}$, $v_{i}(\tilde{M})\geq|\tilde{N}|$.
@@
as desired. It follows that all the conditions required in order to apply Lemma 3 are fulfilled, and we can satisfy agents in $\tilde{N}$ using $|\tilde{N}|-1\leq|N|-2\leq n-2$ distinct copies (recall that at least one agent was already allocated a bundle before). Since $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ is the only place where we produce copies, this concludes the proof. ∎
@@
One can consider a similar question in the context of chore allocation, and ask “what is the maximum number of chores that can be assigned to agents such that no agent’s cost is larger than their maximin share”; that is, how many chores are needed to be set aside in order to produce a full MMS allocation. It turns out that an analogous process to $\mathsf{Match\mbox{-}n\mbox{-}Fill}$ in the context of chores outputs an allocation where at most $n-2$ items are disposed of, and each agent is allocated a bundle of cost no larger than their MMS value. More on this is in Appendix A.
@@ 4.2 Full MMS with $\lfloor m/3\rfloor(1+o(1))$ Copies
$=$ $\left\lfloor\left(\alpha+\frac{1+\alpha}{\frac{1}{1+\alpha}n-1}\right)m\right\rfloor$
@@
5 Approximate-MMS with Copies: Additive Valuations
@@
In this section we study approximate MMS allocations while allowing duplication of goods.
First, in Section 5.1, we convert a fair division instance into an ordered instance with $n$ agents and $m$ items, in which all agents have the same ordinal preferences.
It has been observed, for the original MMS problem (without copies), that it is without loss of generality to consider ordered instances [16, 14]. Notably, where copies are allowed, the same transformation as in [16, 14] may no longer yield a valid allocation, see example in Appendix F. To address this, we impose an additional constraint on the allocation in the ordered instance (see Definition 5), and show that under this constraint, it suffices to consider ordered instances. We then devise algorithms that produce allocations adhering to this constraint.